If x varies directly as​ y, and x=12 when y=2​, find x when y=10.
Answer by Alan3354(69443)

If x varies directly as​ y, and x=12 when y=2​, find x when y=10.
===============================================
Do it like this:
----
If z varies inversely as​ w, and z=30 when w=0.4​, find z when w=10.
------------
z = k/w --- k is a constant
30 = k/0.4
k = 12
==================
find z when w=10.
z = 12/10
z = 1.2